OpenStudy (anonymous):

Solve each inequality. Graph the solution set and write the answer in interval notation. Terrible at math please walk me through the steps 4 less than or equal to k+11 over 4 less then or equal to 4

OpenStudy (anonymous):

\[4\le \frac{ k+11 }{ 4 }\le4\]

OpenStudy (agent0smith):

You can either break the inequality up into two parts, or keep it as is, depending which way you find easier. If you keep it as is, start by multiplying everything by 4, and then subtract 11 from everything.

OpenStudy (agent0smith):

\[\large 4 \times 4\le 4 \times \frac{( k+11)}{ 4 }\le4 \times 4\]

OpenStudy (anonymous):

\[16\le11K \le16\]

OpenStudy (agent0smith):

k+11 is not equal to 11k, you have to keep it as k+11. Also are you sure this is correct \[4\le \frac{ k+11 }{ 4 }\le4 \] seems like you wouldn't have a 4 on both the left and right. Or else it just means \[\frac{ k+11 }{ 4 }\ = 4\]